### v1

# Authors Should Annotate

Abstract
The status quo for labeling text is third-party annotation, but there are many cases where information directly from the document’s source would be preferable over a third-person proxy, especially for egocentric features like sentiment and belief. We introduce author labeling, an annotation technique where the writer of the document itself annotates the data at the moment of creation. We collaborate with a commercial chatbot with over 10,000 users to deploy an author labeling annotation system for subjective features related to product recommendation. This system identifies task-relevant queries, generates on-the-fly labeling questions, and records authors’ answers in real time. We train and deploy an online-learning model architecture for product recommendation that continuously improves from author labeling and find it achieved a 534% increase in click-through rate compared to an industry advertising baseline running concurrently. We then compare the quality and practicality of author labeling to three traditional annotation approaches for sentiment analysis and find author labeling to be higher quality, faster to acquire, and cheaper. These findings reinforce existing literature that annotations, especially for egocentric and subjective beliefs, are significantly higher quality when labeled by the author rather than a third party. To facilitate broader scientific adoption, we release an author labeling service for the research community at academic.echollm.io.
Authors Should Annotate
Marcus Ma∗1,2   Cole Johnson∗1   Nolan Bridges∗1 Jackson Trager2Georgios Chochlakis2 Shrikanth Narayanan2 1 echo LLM  2 University of Southern California {marcus,cole,nolan}@echollm.io equal contribution
1 Introduction
The ultimate goal of language understanding is decoding the underlying intent behind a speaker’s words. Intent cannot be intuited from text alone; true understanding requires contextualizing the message with the source view shaped by the communicator’s individual preferences, beliefs, and judgments. However, when training language models to infer this context, access to the original source is often difficult or impossible. The prevailing method for interpreting source views is via observational proxy, whereby third-party annotators assign labels to text data based on their inferences of what the source author would have believed pustejovsky_2012.
Third-party observers, however, are noisy proxies of the original source. Annotators lack “privileged access” to authors’ authentic internal states vazire_2010 and lack context of the original situation jones_1971; wallace_2014. Decades of annotation research have documented fundamental limitations of third-party annotation, such as annotator biases, misalignment with the author intent, and pervasive annotator disagreement (snow_2008; fort_2016; aroyo_2015; uma_2021). While some of these limitations can be mitigated — for example, in subjective tasks, annotator disagreement may be indicative of multiple interpretations rather than noise chochlakis_2025, and standardized techniques can lead to high inter-annotator agreement paletz_2024 — they are structurally tied to the methodology of third-party annotation. The reliance on third-party annotation is simply a logistical necessity, as it is impossible to reconstruct the original author’s viewpoint during retroactive text labeling.
In this work, we investigate the technical feasibility of collecting annotation data directly from authors in real time. We present author labeling, a new annotation methodology where we present authors with the opportunity to annotate their documents at their moment of creation. In the real-world setting of chatbot conversations, we deploy a data collection system that uses lightweight LLMs to automatically monitor user messages for relevance to various topics and tasks. When this system identifies task-relevant moments, we create a contextualized annotation task based on the user’s query and present it to the user directly in the chatbot interface for them to answer.
Figure 1: Example of an author-labeling task on echo for the query, “want to switch to a new health insurance provider to cover dental”.
To test author labeling, we introduce Echo, or Experimental Coordination of Human Objectives, as an online-learning model architecture that continuously improves through author labeling by strategically querying authors to reduce uncertainty over a predefined set of features or tasks. We implement Echo for the task of product recommendation and deploy it into a live chatbot service with over 10,000 users simultaneously with an industry-standard banner-ad company’s product. We find that Echo achieves a click-through rate (CTR) of 7.52%, a fivefold increase over the industry baseline’s 1.40% CTR. We also directly compare our author-labeling approach to other annotation methods for the task of sentiment analysis. Using three other annotation sources — LLM-as-an-annotator, crowdsource via Mechanical Turk, and hand-selected expert — we find that author labeling is the most performant, internally consistent and cheapest (besides LLM annotation) of the methods. To support broader scientific adoption, we release a publicly available author labeling service at academic.echollm.io, enabling researchers to build datasets grounded in real-time author intent rather than third-party approximation.
2 Author Labeling
Author labeling is an analogous methodology to ecological momentary assessment (EMA), a common technique in behavioral psychology that uses egocentric, in-the-moment self-reports (shiffman_2008). In §2.1, we highlight three reasons why direct self-reports such as author labeling and EMA are preferable to third-party annotation. In §2.2, we provide an overview of how we conducted author labeling in a chatbot setting, and in §2.3, we discuss limitations of author labeling.
2.1 Motivation for Author Labeling
Intent Grounding.
Language is inherently ambiguous, and a single sentence can contain many different interpretations. Even the most skilled and attentive annotators can at best identify ambiguities and make an educated guess, and only the author of the document understands the true intent. oprea_2020 find that annotator performance on sarcasm detection — an example of an ambiguous linguistic construct — achieves an F1-score of 0.616 compared to authors’ actual intents. Additionally, when constructing passages, authors themselves are fully aware of the background, personal context, and other non-linguistic influences that motivate the writing of the passage in the first place. buechel_2017 find that instructions prompting annotators to explicitly consider the author’s context and viewpoint increases label quality.
Ambiguity Resolution.
In subjective annotation tasks, disagreement between annotators can mean more than noise and can in some cases be a useful signal itself for detecting ambiguity beigman_2009; chochlakis_2025. This ambiguity is simply a residue of the annotation process, though; language can be ambiguous to external annotators, but to the author, the meaning should be more clear.
Abstention as Signal.
Subjective annotation tasks see improved performance when providing annotators with an abstaining label such as “not sure” or “none of the above” uma_2022. This hedges against ambiguous data where assigning a label would force annotators to guess, adding noise to the data. When authors themselves label their data with these choices, an abstention selection is no longer just a noise mitigation strategy but a valuable signal for model calibration and confidence. If authors were to label their own document with an abstention, then it is more likely to mean that there is no correct or applicable label to the document. When annotators label it so, they are simply lacking enough information to make a prediction on the author’s internal beliefs.
2.2 Data Collection
The largest obstacle for collecting data via author labeling is access to a diverse, large data source with an integrated UI platform for author annotations. We partner with echo LLM [^1] https://chat.echollm.io, a free-to-use chatbot platform that provides users with personalized product recommendations. Echo relays LLM inference to third-party LLM providers such as OpenAI and Anthropic and monitors the output stream. When recommendation for a product from echo’s catalogue would benefit the user’s conversation, echo generates a conversation-contextualized and LLM-generated advertisement prompted to be as helpful and complementary to the chatbot response as possible.
We conducted author labeling with the purpose of gathering labeled information on a variety of intent features relevant to product recommendation, such as sentiment, urgency, and desired product features. We set up an automatic trigger that ran on every user query, where after a message was sent, we first passed the contents through a spam and greetings filter. Then, we evaluated the message with a lightweight LLM to see if it a given feature was relevant to the conversation. Some features, such as sentiment, were applicable to almost every query, while others, such as purchase urgency, were only relevant when it was clear the user was exhibiting purchase intent. If enough intent features that we needed data on were deemed relevant to the conversation, we presented users with a popup overlay interface asking them four multiple-choice or single-word free-text questions relevant to the conversation (Figure 1). These tasks are generated on-the-fly based on the conversation, and each question is designed to add new information about one specific aspect of the user’s intent that is not immediately answerable from the conversation itself. We provide additional detail in §3.1.
Figure 2: Diagram of the Echo architecture. Feature models are updated via author labeling tasks generated to minimize feature selector uncertainty. The feature selector is trained on losses from the downstream task.
Ensuring data quality.
We designed our label questions to be as non-intrusive and easy-to-answer as possible, drawing on established survey design principles stating that minimizing cognitive burden reduces the likelihood of low-quality or falsified data krosnick_1991; dillman_2014; this is especially important given that the authors themselves were not presented with detailed instructions typically given to third-party annotators. With this in mind, we designed all of our questions to require either one-click or one-word answers and enforced a 5-second minimum delay after presenting the task so that users would read the question rather than rush through. We also relied on echo’s user fraud detection algorithms and only report users whose data has been verified by their system to be not fraudulent (see §B.2 for more details).
In October and November 2025, a total of 928,078 messages were sent on echo from 22,077 users. We filtered out tasks where the conversation revealed PII or discussed sensitive topics, and we also filtered out users who were spamming echo to exclusively receive tasks rather than asking genuine questions. After these filters, we presented 54,977 four-question surveys and achieved a survey completion rate of 84.3%, collecting 185,548 individual author-labeled datapoints.
2.3 Limitations of Author Labeling
Author labeling does not remove all ambiguity and noise from annotation. Different authors can have different interpretations of the same label, a well-known phenomenon for emotion and affect barrett2017emotions. In self-reports, the phrasing of the question and potential labels influences the final answer significantly schwarz_1999. The timing of data collection also greatly influences quality, regardless of the labeler, and even repeated self-reports from the same individual can have variance. Ephemeral details and post-hoc rationalization often change judgment from the original interpretation hoelzemann_2024, but there is no definitive answer on which answer is most “correct”. It is also well-known that inter-annotator variance adds a large source of variance for model performance and that simply adding or changing annotators can greatly change model behavior geva_2019. Moving labeling to the authors themselves will not solve this issue for general systems, as each author still carries their own personalized viewpoint that can change over time. Additionally, author labeling does not solve demographic representation issues; however, when authors themselves annotate their data, authors’ demographics become a feature to consider rather than a potential source of labeling bias. Observer bias and the Hawthorne effect landsberger1958hawthorne can influence author answers, and some subjective tasks such as hate speech identification require a third party to determine if a statement is harmful. People also individually vary on their level of emotional intelligence salovey_1990 and ability to be aware of and distinguish their emotions barrett_1998; bagby_1994.
3 Echo
To test author labeling, we introduce Echo, or Experimental Coordination of Human Objectives, as an online-learning model architecture that continuously improves through author labeling by strategically querying authors to reduce uncertainty over a predefined set of features or tasks. Echo has three distinct phases: feature generation, feature selection, and the downstream task. Feature generation and feature selection are trained through two separate training loops using author labeling and downstream losses.
Figure 3: Click-through rate per day over the course of the 27-day experiment. The industry baseline was presented concurrently with Echo from days 5 to 12. Daily CTR for Echo and the baseline are shown as bars while Echo’s cumulative CTR is shown as the scatterline.
3.1 Feature Generation
We take a raw input $\mathbf{x}$ and create an ensemble of $n$ differentiable feature models, $\mathcal{M}=\{M_{1},\dots,M_{n}\}$ parameterized by $\theta_{M_{i}}$. Each feature model is designed to predict one specific, distinct feature value (represented as a number, a class, a vector representation, etc.), meaning that we assume a priori a set of all potential features that could be used by the downstream task. For a chatbot environment, for example, we could have $x$ to be the user message and $\mathcal{M}$ to be a collection of generative language models each tasked with identifying one specific linguistic or semantic aspect. We use each model to predict its relevant feature value:
$\hat{f}_{i}=M_{i}(\mathbf{x};\theta_{M_{i}})$ (1)
and combine them into a candidate feature pool vector $\mathbf{\hat{F}}$.
3.2 Feature Selector
The feature selector, with parameters $\theta_{S}$, selects the top $k$ elements from $\mathbf{\hat{F}}$. This emphasizes the design decision that $n>>k$, for example having thousands of possible feature models and only picking the top ten for the downstream task [^2] An alternative approach, if running all feature models on every input is expensive, is to train the feature selector to pick the top-k feature models $\{M_{s_{1}},\dots,M_{s_{k}}\}$ rather than the top-k predicted features $\{\hat{f}_{s_{1}}\dots\hat{f}_{s_{k}}\}$..
The selector assigns a relevancy score $s_{i}\in\mathbb{R}$ and an uncertainty estimate $\sigma_{i}\in\mathbb{R}^{+}$ for each feature.
$\mathbf{s},\bm{\sigma}]=\text{Selector}(\mathbf{\hat{F}};\theta_{S})$ (2)
We re-rank the features by their relevance score $s_{i}$ and pick the top-$k$ highest features, which we pass to the downstream model. The relevancy score $s$ is optimized via the downstream model’s loss while the uncertainty $\sigma$ is updated via the author labeling response.
3.3 Downstream Prediction
We use the features passed from the feature selector in our downstream algorithm, which is configured to take exactly $k$ input features, which has a downstream evaluation and loss. In our implementation of a product recommendation system for example, described in further detail in §4, the downstream recommendation algorithm receives specific user-intent features from the feature selector and produces the contextualized recommendation advertisement, where we use the click-through rate and more direct author labeling (i.e., asking if the recommendation was helpful) as our downstream loss.
3.4 Training
We have two forms of feedback to supervise the training of Echo: the author labeling on specific feature predictions, which is used to train the feature models and the confidence of the feature selector, and the downstream loss, which trains the feature selector’s relevance scores and optionally the downstream algorithm itself.
3.4.1 Author Labeling
After the feature selector identifies the top-$k$ features, we use the uncertainty scores $\bm{\sigma}$ and generate tasks to ask the user what the true value of that feature should be, making sure to include the predicted feature value as one of the options. When the author provides the ground-truth label for that feature, we use the value to update both the feature selector’s uncertainty calculations and the associated feature model’s prediction. We use this loss to update only the specific feature model, not all feature models, as the author labeling task is generated solely for that individual feature. To update the feature selector’s uncertainty, we binarize the uncertainty loss into a 0 or 1 depending on if the label chosen by the author matches the feature model’s predicted label.
3.4.2 Downstream Loss
We use the downstream loss to update the feature selector’s feature selections; crucially, we stop backpropagation at the feature selector level. The downstream loss does not update the feature models themselves, since we do not know if the downstream loss was caused by bad predicted features or bad feature selection.
4 Implementing Echo for Product Recommendation
Figure 4: Day-by-day CTR Monday through Sunday.
We implement a version of Echo designed to recommend a product to the user based on their perceived needs from their conversation. Our feature models are lightweight LLMs prompted to guess the value of one of many features deemed potentially important for whether the user would be open to purchasing a product and also what specific product they would want, such as purchase timing and individual product specifications [^3] We omit the full feature list, prompts, and model details to protect proprietary information from echo LLM.. The feature label space differs per-feature, and we use a combination of binary values, free-form text, and a closed set of predefined labels depending on the feature.
The downstream algorithm is echo’s product recommendation system, which takes in the selected features, matches it to the closest item in echo’s product catalogue, and makes a final decision to display an advertisement rewritten to fit naturally into the conversation. Author labeling tasks are generated via an LLM call using a task generation template combined with feature specifications and definitions. We update our feature model LLMs with author labels through a combination of in-context learning, automated prompt engineering zhou_2023, and supervised finetuning. We also update our downstream algorithm through author labeling, where we ask the user for feedback regarding the product recommendation itself.
4.1 Experimental Setup
We deploy our product recommendation implementation of Echo into a real-world chatbot environment. As the chatbot receives a user query, it passes the message quickly through Echo’s lightweight filter, which classifies messages as “taskable” or not (see Table 6 for examples). If the message is taskable, we pass the user query into the feature models and display author-labeling tasks for the selected features. We display these tasks independently of the downstream product recommender’s decision to display an ad.
Before displaying any ads, we passed taskable messages through only phases 1 and 2 (feature generation and selection) with author annotation in order to train the feature models. We experimented with user-specific feature models but found limited success for individidual users with a lot of feedback, and decided ultimately on training feature models on all users’ data to be more generalizable. We trained on about 50,000 author labels for about two weeks before launching the full version of Echo with product recommendation, beginning a 28-day experiment where we measured the click-through rate (CTR) of Echo compared to an industry baseline. Click-through rate is the ratio of user clicks to impressions, a standard measure in the advertisement industry for the effectiveness of our product recommendations.
Industry Baseline
During the four-week experiment, we also displayed advertisements provided by an established advertising company for an 8-day period. This advertiser presented banner ads, which are generated dynamically based on the user’s query via their proprietary keyword-matching algorithm. During this period, we passed user queries to the industry baseline, and if it decided to service an ad, we displayed their advertisement in the same user interface as Echo’s recommendations.
4.2 Results
We plot the day-by-day CTR of Echo and the industry baseline in Figure 3. In total, Echo created 10,378 impressions with 789 clicks for an overall CTR of 7.52%, while the industry baseline presented 1,711 impressions with 24 clicks for a CTR of 1.40%, meaning that Echo achieves a 537% performance increase compared to the industry standard.
We also see a novelty bonus for Echo’s CTR: CTR spikes early into the introduction of each model then drops to a stable rate. This is consistent with advertisement literature finding that novelty perception for new stimuli decays rapidly and that increased exposure to a stimulus reduces advertisement engagement over time Wu_2007; Havlena_2004.
While click-through rates for ad displays in chatbot environments have been understudied, the global average CTR for display ads on Google Display Network (i.e., on internet search) ranges between 0.46% (irvine_2024) and 0.57% (chaffey_2024), with static banner ads seeing CTR as low as 0.10% bannerflow_2024. We believe that the increased performance of the industry baseline, achieving 1.40% CTR in our chatbot environment, could be due to higher levels of attention and novelty for users using chatbots versus internet search.
Day-by-day CTR variations.
We display day-by-day CTR in Figure 4. We observe that Mondays and Fridays have the lowest average CTR with peaks mid-week and on the weekend, aligning with established consumer patterns. Monday’s low performance aligns with the task-orientation online browsing hypothesis Bussiere_2016, stating that consumers allocate their cognitive resources to starting the work week, so buying products is not on their mind. Mid-week jumps could be explained by “cyberloafing” Lim_2002, where online browsing increases as a distraction from work; the Friday dip could be best explained by higher rates of socialization, which is also evidenced by a dip in impressions on Friday (see Figure 9 in Appendix C), with a weekend surge in CTR possibly explained by increased leisure and exploration from online consumers Moe_2004.
Figure 5: Top-ten verticals by CTR with a minimum of 180 impressions.
5 Comparing Author Labeling to 3rd-Party Annotation
We now perform a head-to-head analysis of author labeling for annotation compared to three other annotation methods for the task of sentiment analysis. In the chatbot platform, we introduced an automatic task generator that created multiple-choice questions asking the user about their current emotional state based on their conversation. Both the question and the answers were derived from an automatic LLM generation prompted to create a contextualized question aimed at understanding the user’s emotional state and to come up with several plausible emotions or feelings as options. We intentionally chose to have the LLM come up with labels rather than select options from a predefined emotion set so that the labels themselves could incorporate context, thus further disambiguating the label value to users (i.e. the label “Concerned about pollution” vs. “Concerned”). We filtered out spam and non-English messages, messages from users detected to be bots, and messages with little to no emotional subtext. We also excluded all tasks with any mention of personally identifiable information, religion, NSFW content, non-English, or political content. We restricted our sample to users who had completed a minimum of ten annotation tasks and sampled exactly 6 tasks for every user, and performed a final manual inspection of all user messages and tasks. This yielded 1,962 total tasks across 327 users and 1043 conversations, with each conversation having an average of 1.88 tasks and each user having an average of 3.19 conversations.
5.1 3rd-Party Annotation Sources
In addition to author labeling, we collect three types of annotation: LLM-as-an-annotator, crowdsourcing via Mechanical Turk (MTurk), and expert annotation. We chose these three methods to represent the status quo in annotation covering the high-volume scaling capacity of crowdsourcing snow_2008, the emerging trends of using LLMs as low-cost annotators ding_2023, and the quality control of expert annotation.
We report annotator demographics in Table 5; however, the authors have reason to believe that some of the demographics reported by Mechanical Turk annotators may have been falsified, as the authors received 15 emails from MTurk annotators with traditionally Indian or South Asian names written in non-fluent English, yet only two MTurk annotators reported having an Asian ethnicity.
For the expert annotators, we selected a sample of 300 sentiment tasks from the 1,962 total tasks. We collected three annotations per datum for all annotation methods, totaling 196 Mechanical Turk annotators and 30 expert annotators. For the LLM annotator, we used the same prompt for Llama 3.1 8B, Llama 3.3 70B, and GPT-OSS 20B as the three “annotators”. We present additional annotator details in Appendix B.
Author Accuracy Annotator Agreement KL Divergence LLM 42.76% 0.36 0.01 MTurk 33.18% 0.07 0.12 Expert 42.33% 0.12 0.08
Table 1: Author accuracy, annotator agreement (Cohen’s Kappa), and KL Divergence with the author label distribution for the three annotation methods.
5.2 Alignment with Author Labels
We present overall accuracy with author labels, annotator agreement, and KL divergence with author label distributions for the three annotation methods in Table 1444We note that random accuracy is 25%, as all data are four-choice multiple-choice questions.. We find that experts (42.76%) and LLMs (42.33%) achieve essentially the same level of accuracy, but LLM annotators had much higher agreement ($\kappa$=0.36) than experts (0.12). However, we find that when experts do agree, their overall accuracy is higher (Table 2), achieving the highest accuracy among methods of 63.0% when all annotators agree on a label. Meanwhile, MTurk annotations perform the worst on all three metrics, which is to be expected compared to expert annotators but surprising in comparison to LLM annotation.
5.3 Internal Consistency and Predictivity
Alignment with author labels holds no meaning if the author labels themselves do not carry any useful signal. We conduct an experiment to see how internally consistent and predictive different annotation types are, including author labeling. In this setup, we prompt an LLM to predict what label the user would assign a question with a similar prompt as the LLM annotator’s (Appendix C.2); however, we also provide as context five other questions asked to the same user along with labels from five different sources: the author label, the three 3rd-party annotation methods, and random labels. Our intuition is that the predictor will perform better with the additional information given from other questions, and that the higher quality the label source, the better the performance.
Three Agreements Two Agreements No Agreements LLM 50.5% ($n=746$) 39.6% ($n=1027$) 25.4% ($n=189$) MTurk 37.4% ($n=262$) 35.1% ($n=1223$) 29.1% ($n=477$) Expert 63.0% ($n=46$) 41.7% ($n=175$) 25.3% ($n=79$)
Table 2: Accuracy per annotation method based on number of agreements per datum. For “No Agreements”, one of the three different answers is randomly selected.
Results
We present results in Table 3. Performance follows the same pattern as in Table 1: author labeling yields the highest accuracy, followed by LLM and expert annotators with comparable performance, and MTurk with the lowest overall accuracy. Because only the actual labels themselves changed between different prompts, we assert that the user labels are the most internally consistent and predictive of all methods tested.
We also explored whether it mattered if the task examples occurred in the same conversation in Figure 6. We find that some methods benefit from a higher number of in-conversation task answers while others do not, but also note that this differs by model as well (See Appendix C.3. However, across all models, author labeling benefits the most from additional in-conversation examples, as intuitively additional context from the examples is likely most relevant in the same conversation and timespan of the actual question and author labeling contains the highest fidelity of annotator answer.
Model User Labels LLM Annotator Expert Annotator MTurk Annotator Llama3.1 8B Instruct 43.1±0.44 39.7±0.21 41.6±2.71 37.5±0.80 GPT-OSS 20B 46.2±0.13 41.5±0.16 40.4±3.34 40.2±1.13 Llama3.3 70B Instruct 46.9±0.52 42.7±0.68 43.6±2.46 40.8±0.54
[…]
Figure 6: Predictor accuracy vs. number of in-conversation examples. All prompts always had 5 examples in total.
5.4 Cost Analysis of Annotation Methods
We conduct a simple cost and time analysis comparing author labeling to the two human annotation methods, Mechanical Turk crowdsourcing and expert annotation. We rewarded chatbot users with \$0.08 per task response, with each task taking less than ten seconds to complete (five seconds of forced response delay plus five seconds to answer). This yields a rate of \$28.80 per hour for author labeling tasks. For experts, while they graciously volunteered for free, we consider a rate of \$20 per hour to be reasonable. For Mechanical Turk, we offered a fixed rate of \$5 for a 30-question survey, which we initially estimated to take 30 minutes. In actuality, it took Mechanical Turk annotators an average of 9.1 minutes to complete the survey and expert annotators 17.0 minutes [^5] The 9.1 minute survey completion for MTurk likely indicates these annotators did not entirely read through each conversation, as assuming a fast reading rate of 250 WPM, this would mean they read 50 words per conversation, well below the 125-140 word average.. We present the cost per datum and time per datum in Table 4, which shows author labeling as significantly cheaper and faster than the other annotation methods. We note that our calculations do not factor in that 3rd-party annotation methods typically require multiple annotators per datum; these rates would double or triple when considering additional annotators.
Cost per Datum Time per Datum
Author \$0.08 10 sec
MTurk \$0.17 18.2 sec
Expert \$0.19 34 sec
Table 4: Time and cost analysis of author labeling, Mechanical Turk crowdsourcing, and expert annotation.
6 Conclusion
Reliance on third-party annotation proxies for labeled text data arose out of technical necessity rather than quality improvement. In this work, we demonstrate a technical implementation of author labeling and ask document authors in real time to label their own data. As our two experiments show, author labeling is more internally consistent, cheaper, and faster than traditional 3rd-party annotation techniques and better retains the original author’s signal and intent. We introduce a public author labeling service for the research community at academic.echollm.io to encourage adoption of author labeling as the new gold standard in annotation.
7 Ethics
In order to protect users’ privacy, we will not release any of the personal conversations used in our analysis. We also redact names, locations, and other PII when presenting it to our annotators and omit any conversations dealing with sensitive topics such as personal health or finances. Users fully consent to their data being analyzed for business and scientific purposes when signing up to echo LLM and are fully informed on how they will be compensated for filing out label questions. Filling out label questions is always completely optional and their decision to fill out or not fill out a task does not affect their future experience on echo in any way.
8 Limitations
We previously discussed the theoretical limitations of author labeling in §2.3, but now discuss additional experimental limitations. Because of the real-world nature of our experiments, we could not control for all factors that occurred in the chatbot. The userbase itself was growing throughout the four-week trial period, and some spammy users were banned or excluded from analysis after the fact. Additionally, our head-to-head analysis of 3rd-party annotation to author labeling dealt with egocentric sentiment analysis. We did not analyze objective annotations or non-egocentric subjective tasks, both of which are tasks where being the author of the document might not be as important to annotation. In these scenarios, the qualifications of the author matter, which were factors that were outside the scope of this work. We also used LLMs to generate the questions in order to be as contextually relevant as possible, but the phrasing of questions can greatly influence self-reports schwarz_1999. Additionally, although we introduce the general architectural framework of Echo, we only test its efficacy for the downstream task of product recommendation.
Appendix A Additional Details on Product Recommendation Experiment
A.1 Impression Calculation
We count each unique, individual instance of new content as a single impression, regardless of how many times it was displayed to the user (i.e., reloading when logging back in does NOT increase impression count). A single impression, if clicked multiple times, counts as multiple clicks, which is standard practice in the advertising industry. We noticed that some users would occasionally spam our chatbot with specific phrases with the hope of eliciting task rewards; we banned and excluded flagrant users from analysis, and to counteract additional spammy behavior, we merged impressions that occurred from messages that occurred too quickly in succession, as the only explanation for this user behavior would be for task spamming. We applied this calculation logic to both Echo and the industry baseline.
A.2 Click and Impression Plots
We present impression and click counts instead of CTR for figures 3, 4, and 5 on the next pages.
Figure 7: Impressions per day over the course of the 27-day experiment. The industry baseline was presented concurrently with Echo from days 5 to 12. Daily impressions for Echo and the baseline are shown as bars while Echo’s cumulative impressions are shown as the scatterline.
Figure 8: Clicks per day over the course of the 27-day experiment. The industry baseline was presented concurrently with Echo from days 5 to 12. Daily clicks for Echo and the baseline are shown as bars while Echo’s cumulative clicks are shown as the scatterline.
Figure 9: Day-by-day impressions Monday through Sunday.
Figure 10: Day-by-day clicks Monday through Sunday.
Figure 11: Top-ten verticals by impressions.
Figure 12: Top-ten verticals by clicks.
Appendix B Annotation Details
B.1 Qualtrics Survey
We conduct all annotations via Qualtrics. Participants are informed that their participation is voluntary, confidential, and will be compensated (except for experts who graciously volunteered for free). The survey presented each task on its own page, first displaying the entire user conversation, then the question, then the answers. Long LLM responses were truncated to the first 300 characters while user messages were always retained in full. As a tradeoff between speed and complete contextual fidelity, we only displayed the five most recent messages of the conversation, which the authors acknowledge could in theory limit the total context of the conversation for very long conversations; however in practice, the vast majority of conversation context was unambiguously clear given the previous five messages.
B.2 Non-US and Spam User Filtering
We filtered out several thousand users from analysis based on the following criteria:
  1. 1.
    Location based outside the US
  2. 2.
    Location inside the US, but using a known VPN location inside the US
  3. 3.
    Location inside the US, but individual locations traversing the country several times a day, indicating VPN usage
  4. 4.
    Consistent pattern of messages sent in order to just elicit tasks and their rewards
  5. 5.
    Consistent pattern of messages that were spam or LLM-generated
B.3 Mechanical Turk Details
We filter Mechanical Turk annotators to have over 95% HIT approval rate and to have completed a minimum of 5000 HITs. We create 66 separate batches to annotate the 1,962 datapoints and have three unique annotators per datum. Annotation occurred entirely in November 2025.
B.4 Expert Details
We found expert annotators through snowball sampling. The authors had individual conversations with each expert on instructions for filling out the survey to ensure quality. The authors also personally verified all of the demographic information of each annotator. Annotation occurred from November 2025 to December 2025.
B.5 Annotator Demographics
Characteristic MTurk (N=196) Expert (N=30)
Age
   Mean (SD) 40.2 (10.3) 27.8 (11.4)
   Range 22–73 22–61
Gender, n (%)
   Female 90 (45.9) 7 (23.3)
   Male 105 (53.6) 21 (70.0)
   Non-binary / third gender 0 (0.0) 1 (3.3)
   Prefer not to say 0 (0.0) 1 (3.3)
Race/Ethnicity, n (%)
   White 186 (94.9) 10 (33.3)
   Asian 2 (1.0) 12 (40.0)
   Black 7 (3.6) 1 (3.3)
   White/Asian 0 (0.0) 4 (13.3)
   White or European American,Hispanic or Latino/Latinx 0 (0.0) 1 (3.3)
   White or European American,Native Hawaiian or Pacific Islander 0 (0.0) 1 (3.3)
   Hispanic/Latinx 1 (0.5) 1 (3.3)
Education, n (%)
   High school graduate 4 (2.0) 1 (3.3)
   Some college 7 (3.6) 0 (0.0)
   2 year degree 11 (5.6) 0 (0.0)
   4 year degree 135 (68.9) 21 (70.0)
   Professional degree 32 (16.3) 2 (6.7)
   Doctorate 5 (2.6) 6 (20.0)
Political Party, n (%)
   Democrat 87 (44.4) 21 (70.0)
   Republican 91 (46.4) 3 (10.0)
   Independent 18 (9.2) 5 (16.7)
Political Ideology, n (%)
   Very liberal 4 (2.0) 3 (10.0)
   Liberal 12 (6.1) 6 (20.0)
   Slightly liberal 5 (2.6) 6 (20.0)
   Moderate 11 (5.6) 9 (30.0)
   Slightly conservative 24 (12.2) 2 (6.7)
   Conservative 108 (55.1) 3 (10.0)
   Very conservative 29 (14.8) 1 (3.3)
Table 5: Demographics of Mechanical Turk and Expert Annotators.
B.6 LLM Annotator Prompt
We used the following prompt to act as our LLM annotator. We used Llama 3.1 8B Instruct, Llama 3.3 70B Instruct, and GPT OSS 20B as our three “annotators”. Text in blue corresponds to the conversation, red corresponds to the question, and green corresponds to the task options that are substituted into each prompt.
Prompt Template for LLM Annotator You are a human subjects annotator tasked with guessing the answer to a sentiment-related multiple choice question from the perspective of a user. You will be provided with the user’s conversation, which you should use to inform your answer. --BEGIN CONVERSATION--
USER QUERY: Will coffee make the hair smell like coffee and for how long until the smell disappears?
CHATBOT RESPONSE: Yes, using coffee as a hair dye can leave a subtle coffee scent in your hair. The smell may linger for a few hours to a couple of days, depending on your hair type and how well you rinse it out. If the scent is strong initially, it should fade as your hair dries and with subsequent washes. Using a m...
USER QUERY: Will coffee dye leave a stain on clothing after it has set into the hair?
--END CONVERSATION--
Now, from the user’s perspective, how would they answer this question? --BEGIN QUESTION--
How do you feel about the potential coffee smell in your hair after using a coffee dye?
--END QUESTION--
--BEGIN OPTIONS--
A. Frustrated
B. A bit annoyed
C. Neutral
D. Wanting to avoid it
--END OPTIONS--
Respond with exactly one character: A, B, C, or D.
Do NOT add explanations, punctuation, quotes, or any other text.
Respond ONLY with one of these four exact strings:
A
B
C
D
Choose one and output ONLY that line, with no additional characters or text.
Appendix C Additional Details on Sentiment Analysis Experiment
C.1 Dataset Details
We present the number of tasks per conversation in Figure 13 and example tasks in Table 6.
User Message Task Question Options
How to make a blanket? What motivated you to want to make a blanket? A. To relax and unwind B. As a gift for someone C. For personal use D. Out of curiosity
Are there any 96 degree indoor pools around Panama City, FL What’s driving your need for a 96-degree pool in Panama City? A. Looking to relax B. Need stress relief C. Just curious D. Seeking comfort
How to start an online business How do you feel about starting your online business? A. Excited and motivated B. Nervous about the risks C. Unsure where to begin D. Determined to succeed
Table 6: Sample of three taskable messges and the sentiment analysis task question generated for it. Conversation context is not shown here but was to annotators.
Figure 13: Distribution of the 1,962 sampled sentiment tasks across conversation.
C.2 Prompt for Task Predictor
We used the following prompt to act as our task predictor. Text in blue corresponds to the conversation, red corresponds to the question, purple corresponds to examples and answers, and green corresponds to the task options that are substituted into each prompt.
Prompt Template for LLM Annotator You are a human subjects annotator tasked with guessing the answer to a sentiment-related multiple choice question from the perspective of a user. You will be provided with the user’s conversation, which you should use to inform your answer. --BEGIN CONVERSATION--
USER QUERY: Will coffee make the hair smell like coffee and for how long until the smell disappears?
CHATBOT RESPONSE: Yes, using coffee as a hair dye can leave a subtle coffee scent in your hair. The smell may linger for a few hours to a couple of days, depending on your hair type and how well you rinse it out. If the scent is strong initially, it should fade as your hair dries and with subsequent washes. Using a m...
USER QUERY: Will coffee dye leave a stain on clothing after it has set into the hair?
--END CONVERSATION--
The user has also filled out other sentiment-related multiple-choice questions. These are answers to questions from the same conversation:
MESSAGE: is it easy to make a DIY natural dye?
QUESTION: How do you feel about the uncertainty of DIY natural dye duration?
ANSWER: Concerned
These are answers to questions from different conversations:
MESSAGE: How do Airbnb stays compare to boutique hotels for long-term travel?
QUESTION: How do you feel about the uncertainty of long-term travel accommodations?
ANSWER: Anxious about the unknown
MESSAGE: no, i want the tiktok styles
QUESTION: What’s the main emotion driving your school meal rant?
ANSWER: Frustration
MESSAGE: tell me more about it
QUESTION: What’s driving your desire to learn more about swimming?
ANSWER: Curiosity
MESSAGE: Is Hulu cheap?
QUESTION: How do you feel about the cost of streaming services like Hulu?
ANSWER: I’m looking for good value
If applicable, you should use inferred or explicit information about the user’s emotional state from these prior answers, and factor in whether they came from the same or different conversation as this question. If they are not applicable or unrelated, then you do NOT need to incorporate their answers in your final decision. Now, from the user’s perspective, how would they answer this question? --BEGIN QUESTION--
How do you feel about the potential coffee smell in your hair after using a coffee dye?
--END QUESTION--
--BEGIN OPTIONS--
A. Frustrated
B. A bit annoyed
C. Neutral
D. Wanting to avoid it
--END OPTIONS--
Respond with exactly one character: A, B, C, or D.
Do NOT add explanations, punctuation, quotes, or any other text.
Respond ONLY with one of these four exact strings:
A
B
C
D
Choose one and output ONLY that line, with no additional characters or text.
C.3 Predictor Accuracy vs. In-Conversation Examples for Additional Models
Figure 14: Predictor accuracy vs. number of in-conversation examples for Llama 3.1 8B Instruct.
Figure 15: Predictor accuracy vs. number of in-conversation examples for GPT OSS 20B.
C.4 Predictor Accuracy vs. Conversation Length
Figure 16: Predictor accuracy vs. conversation length for Llama 3.1 8B Instruct.
Figure 17: Predictor accuracy vs. conversation length for GPT OSS 20B.
Figure 18: Predictor accuracy vs. conversation length for Llama 3.3 70B Instruct.